OpenStudy (jmartinez638):

find the local minimum and maximum points and points of inflection for the function f(x) = 1/3X^3 - x^2 - 35x +17

OpenStudy (jmartinez638):

\[f(x) =\frac{ 1 }{ 3 } x^3 - x^2 - 35x +17\]

OpenStudy (solomonzelman):

What is the first derivative?

OpenStudy (solomonzelman):

if you are allowed to use the graph, you can get the local max and min from there https://www.desmos.com/calculator/tgu0exzdsv

OpenStudy (solomonzelman):

for points of inflection we need second derivatiev f''(x).

OpenStudy (jmartinez638):

F'(x) = x^2 -2x - 35

OpenStudy (solomonzelman):

yes, and the second derivative?

OpenStudy (jmartinez638):

F''(x) = 2x-2

OpenStudy (solomonzelman):

Yes...

OpenStudy (solomonzelman):

Set f''(x)=0, to solve for inflections but there is one more little test to see if inflections are inflections

OpenStudy (solomonzelman):

say you get x=a, b for f''(x)=0. You need to varify that the concavity is indded changing at x=a and x=b. let's take some number c that is sufficiently close to a (won't use epsilon or any of that), AND find find f''(x-c) and f''(x+c). If they have different signs, then the concavity indeed changed at x-a. (Same for x=b)

OpenStudy (solomonzelman):

Alright, let's go together... f''(x)=0 2x-2=0 x=1

OpenStudy (solomonzelman):

So x=1 is your (only) possible inflection point.

OpenStudy (jmartinez638):

Ok. So to see if it is an actual inflection point, I am assuming we have to see if the direction of the function actually changes. How do we do this?

OpenStudy (solomonzelman):

Yes, very good.... and for clarifying the direction change, we use the test I mentioned: ---------------------------- If you get x=a for f''(x)=0. You need to varify that the concavity is indeed changing at x=a. let's take some number c that is sufficiently close to a (won't use epsilon or any of that), AND find find f''(x-c) and f''(x+c). If they have different signs, then the concavity indeed changed at x-a. (Same for x=b)

OpenStudy (solomonzelman):

So we can take c=0.05 f''(1-0.05) and f''(1+0.05) if they have different signs then concaavity changes, and x=1 is indded an inflection.

OpenStudy (solomonzelman):

f''(1-0.05)=f''(0.95)=2(0.95)-2 f''(1+0.05)=f''(1.05)=2(1.05)-2

OpenStudy (solomonzelman):

you can tell that 2(0.95) is less than 2, so the first equation is negative. AND you can tell that 2(1.05) is greater than 2, so the second equation is positive.

OpenStudy (solomonzelman):

So, fill in blank; x=1, is indeed the __________ ____\(\bf.\)

OpenStudy (solomonzelman):

if you have some questions, ask

OpenStudy (jmartinez638):

X = 1 is indeed the inflection point.

OpenStudy (jmartinez638):

Because there is a sign change on either side of x = 1

OpenStudy (jmartinez638):

That makes a lot of sense, thanks for the help.

OpenStudy (solomonzelman):

Yw